1.|3-5i|
(3-5i)-(3+5i)
(3-3)(-5i-5)i
0-10i
For this problem i took the absolute value and made it in two seperate problems. After i did that i used foil, and seperated again problems. AFter that i solved what was in the parenthesis, then i 0-10i for the answer. I dont i did it right i asked my brother he said it was wrong i dont know if it is right.

Answer by venugopalramana(3286)

.|3-5i|
ABSOLUTE VALUE OR MOD OF A COMPLEX NUMBER X+iY IS GIVEN BY SQUAREROOT OF (X^2+Y^2) AND IS SYMBOLISED AS
|X+iY|=SQRT(X^2+Y^2)...HENCE WE GET
|3=5i|=SQRT.(3^2+5^2)=SQRT(9+25)=SQRT(34)